OpenStudy (anonymous):

y = x^2 + x + 1 has A)2 rational x-intercepts B)2 irrational x-intercepts C)1 rational x-intercept D)0 x-intercepts

OpenStudy (anonymous):

for x-intercepts solve x^2 + x + 1 =0 \[x=(-1 \pm \sqrt{1-4})/2=(-1 \pm \sqrt{-3})/2\] so you will have no x-intercepts, becouse roots are complex

jimthompson5910 (jim_thompson5910):

Alternatively... Use the discriminant formula D = b^2 - 4ac D = (1)^2 - 4(1)(1) ... Plug in a = 1, b = 1 and c = 1 D = 1 - 4 D = -3 Since the discriminant is negative, this means that there are no real solutions. So there are no x-intercepts.

OpenStudy (anonymous):

what bout for y = 4x^2 + 12x + 9

OpenStudy (anonymous):

do the same process again

OpenStudy (anonymous):

so i just do B^2-4ac

jimthompson5910 (jim_thompson5910):

Same thing, but now a = 4, b = 12 and c = 9 D = b^2 - 4ac D = (12)^2 - 4(4)(9) D = 144 - 144 D = 0 Even though we followed the same process, the fact that we got D = 0 tells us that there are 2 identical rational (and real) solutions.

OpenStudy (anonymous):

what if its a postive like D=9

jimthompson5910 (jim_thompson5910):

Then you get two distinct real solutions Since 9 is a perfect square, this means that these two solutions are also rational (ie fractions)

OpenStudy (anonymous):

thanks guys :)

jimthompson5910 (jim_thompson5910):

sure thing